# On Speeding Up Language Model Evaluation

### v2

@@ Appendix A Additional Details - Datasets
Figure 5: Bar plot of model performance on all examples in all six datasets. Examples are ordered by their rank. Dataset $H_{1}$ values are displayed next to dataset name. It can be seen that datasets, such as AlpacaEval, that have a large gap between the best and second best method have small $H_{1}$ values. The $H1$ value indicates the difficulty of identifying the best method with smaller values indicating easier identification.
@@
To simulate a prompt engineering use case, we create two datasets: GSM8K Prompts and PIQA Prompts. For these datasets, we ask GPT4 to generate 205 and 177 prompts following the prompt engineering work from Yang et al. [60]. The LLM used to perform inference on the datasets are Mistra-7B [32] and Tulu-7B [59] respectively. We evaluate the prompts on 784 and 1546 questions on the training set (about 10% of the size of the two training sets). For scoring, we extract the final answers from the responses generated by the LLMs and compare them with the ground truth answers (for GSM8K) or ground truth choice (for PIQA).
@@
Table 3: Ratio of the largest to second largest eigenvalue.
Dataset Name GSM8K Prompts PIQA Prompts GSM8K Models PIQA Models AlpacaEval (Drop Annotator) AlpacaEval
$\frac{\lambda_{1}(E)}{\lambda_{2}(E)}$ 6.0729 11.2547 3.0048 5.0712 2.8573 2.7316
Figure 6: Bar plot of the explained variance of each principal component for all six datasets.
Appendix B Additional Details - Algorithms
@@
1:  The observation matrix $O:=\{0\}^{n\times m}$, the observed scoring matrix $E^{\rm obs}:=\{?\}^{n\times m}$.
@@
3:     Select: Draw uniformly at random $(i,j)\in\{[n]\times[m]|O_{ij}=0\}$.
4:     Evaluate: Run inference for the method-example pair $(f_{i},x_{j})$, score the result, and receive $e(f_{i}(x_{j}))$; $E^{\rm obs}_{ij}\leftarrow e(f_{i}(x_{j}))$.
5:     Update: $O_{i,j}\leftarrow 1$;
@@
Return: $\hat{i}^{*}=\arg\max_{i}\frac{\sum_{j=1}^{m}O_{ij}\cdot E^{\rm obs}_{ij}}{\sum_{j=1}^{m}O_{ij}}.$
@@
1:  The observation matrix $O:=\{0\}^{n\times m}$, the observed scoring matrix $E^{\rm obs}:=\{?\}^{n\times m}$, initial random example index $j$.
@@
3:     Select: Draw uniformly at random $i\in\{[n]|O_{ij}=0\}$.
4:     Evaluate: Run inference for the method-example pair $(f_{i},x_{j})$, score the result, and receive $e(f_{i}(x_{j}))$; $E^{\rm obs}_{ij}\leftarrow e(f_{i}(x_{j}))$.
5:     Update: $O_{i,j}\leftarrow 1$; If $\sum_{i=1}^{n}O_{ij}=n$, draw uniformly at random $j\in\{[m]|\sum_{i=1}^{n}O_{ij}=0\}$.
@@
Return: $\hat{i}^{*}=\arg\max_{i}\frac{\sum_{j=1}^{m}O_{ij}\cdot E^{\rm obs}_{ij}}{\sum_{j=1}^{m}O_{ij}}.$
Algorithm 5 LRF ($\mathcal{A}_{lrf}(T,\mathcal{F},\mathcal{X};\mathcal{M},r,S,T_{0})$)
Input: The evaluation budget $T$, a set of methods $\mathcal{F}$, a set of examples $\mathcal{X}$, the low-rank factorization model $\mathcal{M}$, the rank $r$, the ensemble size $S$, the warm-up budget $T_{0}$.
@@
1:  Uniformly draw $T_{0}$ method-example pairs from $[n]\times[m]$ and get the observation matrix $O\in\{0,1\}^{n\times m}$ and observed scoring matrix $E^{\rm obs}\in\left([0,1]\cup\{?\}\right)^{n\times m}$ w.r.t. these $T_{0}$ evaluations.
@@
3:     Select: Draw uniformly at random $(i,j)\in\{[n]\times[m]|O_{ij}=0\}$.
4:     Evaluate: Run inference for the method-example pair $(f_{i},x_{j})$, score the result, and receive $e(f_{i}(x_{j}))$; $E^{\rm obs}_{ij}\leftarrow e(f_{i}(x_{j}))$.
5:     Update: $O_{i,j}\leftarrow 1$; $\hat{E}\leftarrow\mathcal{M}(E^{\rm obs},O;r,S)$.
@@
Return: $\hat{i}^{*}=\arg\max_{i}\frac{1}{m}\sum_{j=1}^{m}(O_{ij}E^{\rm obs}_{ij}+(1-O_{ij})\hat{E}_{ij}).$
Algorithm 6 UCB-E-LRF (Score Only) ($\mathcal{A}_{uel-so}(T,\mathcal{F},\mathcal{X};\mathcal{M},r,S,T_{0},\eta)$)
Input: The evaluation budget $T$, a set of methods $\mathcal{F}$, a set of examples $\mathcal{X}$, the low-rank factorization model $\mathcal{M}$, the rank $r$, the ensemble size $S$, the warm-up budget $T_{0}$, the uncertainty scaling $\eta$.
@@
1:  Uniformly draw $T_{0}$ method-example pairs from $[n]\times[m]$ and get the observation matrix $O\in\{0,1\}^{n\times m}$ and observed scoring matrix $E^{\rm obs}\in\left([0,1]\cup\{?\}\right)^{n\times m}$ w.r.t. these $T_{0}$ evaluations.
2:  $\hat{E},R\leftarrow\mathcal{M}(E^{\rm obs},O;r,S)$; $\forall f_{i}\in\mathcal{F},B_{i}\leftarrow\frac{1}{m}\sum_{j=1}^{m}(O_{ij}E^{\rm obs}_{ij}+(1-O_{ij})\hat{E}_{ij}+\eta R_{ij})$
@@
4:     Select: Draw uniformly at random $i\in\arg\max_{k\,\mid\,(\sum_{j=1}^{m}O_{kj})\neq m}B_{k}$; Draw uniformly at random $j\in\{k\in[m]|O_{ik}=0\}$.
5:     Evaluate: Run inference for the method-example pair $(f_{i},x_{j})$, score the result, and receive $e(f_{i}(x_{j}))$; $E^{\rm obs}_{ij}\leftarrow e(f_{i}(x_{j}))$.
6:     Update: $O_{i,j}\leftarrow 1$; $\hat{E}\leftarrow\mathcal{M}(E^{\rm obs},O;r,S)$; $\forall f_{i}\in\mathcal{F},B_{i}\leftarrow\frac{1}{m}\sum_{j=1}^{m}(O_{ij}E^{\rm obs}_{ij}+(1-O_{ij})\hat{E}_{ij}+\eta R_{ij})$.
@@
Return: $\hat{i}^{*}=\arg\max_{i}\frac{1}{m}\sum_{j=1}^{m}(O_{ij}E^{\rm obs}_{ij}+(1-O_{ij})\hat{E}_{ij}).$

### v3

@@ Appendix C Additional Details - Datasets
Figure 5: Bar plot of model performance on all examples in all six datasets. Examples are ordered by their rank. Dataset $H_{1}$ values are displayed next to dataset name. It can be seen that datasets, such as AlpacaEval, that have a large gap between the best and second best method have small $H_{1}$ values. The $H_{1}$ value indicates the difficulty of identifying the best method with smaller values indicating easier identification.
@@
To simulate a prompt engineering use case, we create two datasets: GSM8K Prompts and PIQA Prompts. For these datasets, we ask GPT4 to generate 205 and 177 prompts following the prompt engineering work from Yang et al. [2023]. The LLM used to perform inference on the datasets are Mistra-7B Jiang et al. [2023] and Tulu-7B Wang et al. [2023] respectively. We evaluate the prompts on 784 and 1546 questions on the training set (about 10% of the size of the two training sets). For scoring, we extract the final answers from the responses generated by the LLMs and compare them with the ground truth answers (for GSM8K) or ground truth choice (for PIQA).
@@
Table 3: Ratio of $i^{th}$ singular value to the largest singular value of $S$ in each setting. $\sigma_{i}$ represents the $i^{th}$ largest singular value of $S$. The ratio $\frac{\sigma_{r+1}}{\sigma_{1}}$ can be interpreted as the relative reconstruction error with the best rank-$r$ approximation of $S$. For many datasets, the small numerical differences between the first row and subsequent rows suggest that using rank-1 approximation for UCB-E-LRF is appropriate.
Singular Value Ratio GSM8K Prompts PIQA Prompts GSM8K Models PIQA Models AlpacaEval (Drop Annotator) AlpacaEval
$\sigma_{2}/\sigma_{1}$ 0.1647 0.0889 0.3328 0.1972 0.3500 0.3661
$\sigma_{3}/\sigma_{1}$ 0.1588 0.0763 0.2667 0.1758 0.2153 0.2120
$\sigma_{4}/\sigma_{1}$ 0.1538 0.0739 0.2611 0.1715 0.1746 0.1754
Figure 6: Bar plot of the explained variance of each principal component for all six datasets. The first principal component is much more pronounced than other principal components, suggesting the dataset matrices are generally low-rank.
Appendix D Additional Details - Algorithms
@@
1:  The observation matrix $O:=\{0\}^{m\times n}$, the observed scoring matrix $S^{\rm obs}:=\{?\}^{m\times n}$.
@@
3:     Select: Draw uniformly at random $(i,j)\in\{[m]\times[n]|O_{ij}=0\}$.
4:     Evaluate: Run inference for the method-example pair $(f_{i},x_{j})$, score the result, and receive $s(f_{i}(x_{j}))$; $S^{\rm obs}_{ij}\leftarrow s(f_{i}(x_{j}))$.
5:     Update: $O_{ij}\leftarrow 1$;
@@
Return: $\hat{i}^{*}=\arg\max_{i}\frac{\sum_{j=1}^{n}O_{ij}\cdot S^{\rm obs}_{ij}}{\sum_{j=1}^{n}O_{ij}}.$
@@
1:  The observation matrix $O:=\{0\}^{m\times n}$, the observed scoring matrix $S^{\rm obs}:=\{?\}^{m\times n}$, initial random example index $j$.
@@
3:     Select: Draw uniformly at random $i\in\{[m]|O_{ij}=0\}$.
4:     Evaluate: Run inference for the method-example pair $(f_{i},x_{j})$, score the result, and receive $s(f_{i}(x_{j}))$; $S^{\rm obs}_{ij}\leftarrow s(f_{i}(x_{j}))$.
5:     Update: $O_{ij}\leftarrow 1$; If $\sum_{i=1}^{m}O_{ij}=m$, draw uniformly at random $j\in\{[n]|\sum_{i=1}^{m}O_{ij}=0\}$.
@@
Return: $\hat{i}^{*}=\arg\max_{i}\frac{\sum_{j=1}^{n}O_{ij}\cdot S^{\rm obs}_{ij}}{\sum_{j=1}^{n}O_{ij}}.$
Algorithm 5 LRF ($\mathcal{A}_{lrf}(T,\mathcal{F},\mathcal{X};\mathcal{M},r,C,T_{0})$)
Input: The evaluation budget $T$, a set of methods $\mathcal{F}$, a set of examples $\mathcal{X}$, the low-rank factorization model $\mathcal{M}$, the rank $r$, the ensemble size $C$, the warm-up budget $T_{0}$.
@@
1:  Uniformly draw $T_{0}$ method-example pairs from $[m]\times[n]$ and get the observation matrix $O\in\{0,1\}^{m\times n}$ and observed scoring matrix $S^{\rm obs}\in\left([0,1]\cup\{?\}\right)^{m\times n}$ w.r.t. these $T_{0}$ evaluations.
@@
3:     Select: Draw uniformly at random $(i,j)\in\{[m]\times[n]|O_{ij}=0\}$.
4:     Evaluate: Run inference for the method-example pair $(f_{i},x_{j})$, score the result, and receive $s(f_{i}(x_{j}))$; $S^{\rm obs}_{ij}\leftarrow s(f_{i}(x_{j}))$.
5:     Update: $O_{ij}\leftarrow 1$; $\hat{S}\leftarrow\mathcal{M}(S^{\rm obs},O;r,C)$.
@@
Return: $\hat{i}^{*}=\arg\max_{i}\frac{1}{n}\sum_{j=1}^{n}(O_{ij}S^{\rm obs}_{ij}+(1-O_{ij})\hat{S}_{ij}).$
Algorithm 6 UCB-E-LRF (Score Only) ($\mathcal{A}_{uel-so}(T,\mathcal{F},\mathcal{X};\mathcal{M},r,C,T_{0},\eta)$)
Input: The evaluation budget $T$, a set of methods $\mathcal{F}$, a set of examples $\mathcal{X}$, the low-rank factorization model $\mathcal{M}$, the rank $r$, the ensemble size $C$, the warm-up budget $T_{0}$, the uncertainty scaling $\eta$.
@@
1:  Uniformly draw $T_{0}$ method-example pairs from $[m]\times[n]$ and get the observation matrix $O\in\{0,1\}^{m\times n}$ and observed scoring matrix $S^{\rm obs}\in\left([0,1]\cup\{?\}\right)^{m\times n}$ w.r.t. these $T_{0}$ evaluations.
2:  $\hat{S},R\leftarrow\mathcal{M}(S^{\rm obs},O;r,C)$; $\forall f_{i}\in\mathcal{F},B_{i}\leftarrow\frac{1}{n}\sum_{j=1}^{n}(O_{ij}S^{\rm obs}_{ij}+(1-O_{ij})\hat{S}_{ij}+\eta R_{ij})$
@@
4:     Select: Draw $i\in\arg\max_{k\,\mid\,(\sum_{j=1}^{n}O_{kj})\neq n}B_{k}$; Draw uniformly at random $j\in\{k\in[n]|O_{ik}=0\}$.
5:     Evaluate: Run inference for the method-example pair $(f_{i},x_{j})$, score the result, and receive $s(f_{i}(x_{j}))$; $S^{\rm obs}_{ij}\leftarrow s(f_{i}(x_{j}))$.
6:     Update: $O_{ij}\leftarrow 1$; $\hat{S},R\leftarrow\mathcal{M}(S^{\rm obs},O;r,C)$; $\forall f_{i}\in\mathcal{F},B_{i}\leftarrow\frac{1}{n}\sum_{j=1}^{n}(O_{ij}S^{\rm obs}_{ij}+(1-O_{ij})\hat{S}_{ij}+\eta R_{ij})$.
@@
Return: $\hat{i}^{*}=\arg\max_{i}\frac{1}{n}\sum_{j=1}^{n}(O_{ij}S^{\rm obs}_{ij}+(1-O_{ij})\hat{S}_{ij}).$
Appendix E Table of Notations
The notations used in the paper are described below.
Table 4: Notations used in the paper.
Symbol Description
=$\mathcal{F}=\{f_{1},\ldots,f_{m}\}$ A list of LLM-based methods
$\mathcal{X}=\{x_{1},\ldots,x_{n}\}$ A dataset of examples
$s$ A scoring function
$S\in\left([0,1]\cup\{?\}\right)^{m\times n}$ A scoring matrix computed as $S_{ij}:=s(f_{i}(x_{j}))$
$\mu_{i}$ The average score of a method $f_{i}$ on all examples as $\frac{1}{n}\sum_{j=1}^{n}S_{ij}$
$\mu_{i}^{*}$ The average score of the best method
$T$ An evaluation budget with $T$ method-example queries
${\mathcal{A}}(T,\mathcal{F},\mathcal{X})$ An algorithm that predicts the best method among $\mathcal{F}$ on $\mathcal{X}$ with $T$ query budget
$S^{\rm obs}\in\left([0,1]\cup\{?\}\right)^{m\times n}$ A partially observed scoring matrix
$O\in\{0,1\}^{m\times n}$ An observation matrix where 1 means the corresponding entry in $S$ has been observed
$B_{i}$ An upper confidence bound for a method $f_{i}$
$a$ A hyperparameter for UCB-E algorithm that balances exploration and exploitation
$H_{1}=\sum_{i=1,i\neq i^{*}}^{m}\frac{1}{(\mu_{i}-\mu_{i}^{*})^{2}}$ The dataset hardness quantity
$U\in\mathbb{R}^{m\times r}$, $V\in\mathbb{R}^{n\times r}$ The low-rank factorization of $S^{\rm obs}$ such that $S^{\rm obs}\approx UV^{\top}$
$\hat{S}\in\mathbb{R}^{m\times n}$ An estimated scoring matrix $\hat{S}=UV^{T}$ for UCB-E-LRF
$\hat{\mu}_{i}$ An estimated average score of a method $f_{i}$ for UCB-E-LRF
$C$ The ensemble size for UCB-E-LRF
$R\in\mathbb{R}^{m\times n}$ The uncertainty matrix for UCB-E-LRF
$T_{0}$ The warmup budget for UCB-E-LRF
$r$ The rank of low-rank factorization
$\mathcal{M}(S^{\rm obs},O;r,C)$ The low-rank factorization model that outputs $\hat{S}$ and $R$ for UCB-E-LRF
$\eta$ The uncertainty scaling factor for UCB-E-LRF